OpenStudy (anonymous):

Im trying to teach my nephew how to get the common factor from 8, 4, 10. It is easy for me but I dont really know how to help her. Please teach me how to do that.

OpenStudy (anonymous):

write out all factors look for same ones in all of them 1, 2, 4, 8 1, 2, 3, 1, 2, 4, 10 common factors: 1, 2

OpenStudy (anonymous):

ahhh eassy. Thanks again

OpenStudy (anonymous):

cheers ^_^

OpenStudy (radar):

The common factor for those three numbers other than 1, is 2. 2 is common to all three

OpenStudy (anonymous):

I'm pretty sure i wrote that? :x

OpenStudy (radar):

Well what confused me was you had listed 3 as a factor, all be it not common, but I don't see 3 as a factor for any of those numbers.

OpenStudy (anonymous):

oooo so sorry x_X EMUN i made a couple of typos look for same ones in all of them 1, 2, 4, 8 1, 2, 4 <-- here 1, 2, 5, 10 ^-- and here thanks radar

OpenStudy (anonymous):

but it's still 1, 2 i got lucky :P

OpenStudy (anonymous):

Can you help me with this 1/2, 3,2, 1/2. I know the common term is 1/2. But how to teach it in a simple way?

OpenStudy (radar):

You were right, in fact I missed the last line, guess I was focused on the 3 lol

OpenStudy (anonymous):

I don't think I've done common factors with fractions... actually, i was just doing it, i dont' think this is possible....whats your take radar? there are unlimited fraction pairs that can be used to obtain 1/2, 2, 3...

OpenStudy (radar):

I agree, using fractions as factors is not normally done, usually integers when dealing with numbers.

OpenStudy (anonymous):

Im not sure of this, but i do the following: Put the fraction in its simplest terms. 1/2=1*1/2 3/2=3*1/2 1/2=1*1/2 The common term is 1/2. This is the way my brain processes the problem. Again, im not sure of it. Lets try with another example. You give me one please

OpenStudy (radar):

well in that case what is to prevent you from saying 1/2=2* 1/4, 4*1/8 etc.

OpenStudy (anonymous):

right, but i can argue that 1/2 = 4/3 * 3/8 3/2 = 4/3 * 1/2 1/2 = 4/3 * 3/8 then 4/3 is also a common factor you have INFINITY pairs with fractions i think this question is invalid unless the answer is that there ISN'T one? Oo

OpenStudy (anonymous):

ahhh OK. Got it. YOu mean there is not a way of finding a common factor from a fraction?

OpenStudy (anonymous):

I don't believe so.

OpenStudy (radar):

What are you studying, addition and subtracting of fractions?

OpenStudy (anonymous):

But there is not an specific rule for doing this. We have to think about it first

OpenStudy (radar):

Of course fractions have factors, but there can be lots and lots of them, you must have some purpose when factoring fractions?

OpenStudy (anonymous):

the specific rule for finding common factors is first writing out all factors for each number and then seeing the common ones however, like i said before, with fractions, you have an unlimited amount of factors, you can find an unlimited amount of common factors

OpenStudy (anonymous):

Ok thanks